OpenStudy (anonymous):

1.State the vertical asymptote(s), x - intercept(s), y-intercept(s), domain and range of: f(x) =1/4x^2 - 1

OpenStudy (anonymous):

x cannot be 0 and y cannot be negative 1

OpenStudy (anonymous):

there is no y intercept..

OpenStudy (anonymous):

x intercept is when y = 0 1/4x^2=1 4x^2=1 x^2=1/4 x= +1/2 and -1/2